OpenStudy (anonymous):

A rectangular prism has a volume of 360in^3. It has a sqaure base of x centimeters and a height of x + 4. What are the dimensions.

OpenStudy (mathstudent55):

How do you find the volume of a rectangular prism?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

so the base is in centimeters?

OpenStudy (anonymous):

V = 360 in^3 or 5899.34 cu. cm.

OpenStudy (anonymous):

V = l x w x h 5899.34 cm^3 = l x w x h since x is square, thus, x = l = w... and h = x + 4, 5899.34 = x * x * x+4 5899.34 = x^3 + 4x^2

OpenStudy (anonymous):

x^3 + 4x^2 - 5899.34 = 0 ---> equate that haha